# Revisiting Randomization in Greedy Model Search

### v1

@@ Abstract
Combining randomized estimators in an ensemble, such as via random forests, has become a fundamental technique in modern data science, but can be computationally expensive. Furthermore, the mechanism by which this improves predictive performance is poorly understood. We address these issues in the context of sparse linear regression by proposing and analyzing an ensemble of greedy forward selection estimators that are randomized by feature subsampling—at each iteration, the best feature is selected from within a random subset. We design a novel implementation based on dynamic programming that greatly improves its computational efficiency. Furthermore, we show via careful numerical experiments that our method can outperform popular methods such as lasso and elastic net across a wide range of settings. Next, contrary to prevailing belief that randomized ensembling is analogous to shrinkage, we show via numerical experiments that it can simultaneously reduce training error and degrees of freedom, thereby shifting the entire bias-variance trade-off curve of the base estimator. We prove this fact rigorously in the setting of orthogonal features, in which case, the ensemble estimator rescales the ordinary least squares coefficients with a two-parameter family of logistic weights, thereby enlarging the model search space. These results enhance our understanding of random forests and suggest that implicit regularization in general may have more complicated effects than explicit regularization.
Keywords
Regression |||| Forward Selection |||| Greedy Model Search |||| Random Forests |||| Ensemble Learning
@@ 1 Introduction
Sparse linear regression is a well-studied problem in statistics, having important applications to machine learning, signal processing, and other fields. Given a response vector $\mathbf{y}=(y_{1},y_{2},\dots,y_{n})^{\top}$ and data matrix $\mathbf{X}=\begin{bmatrix}\mathbf{x}_{1}&\mathbf{x}_{2}&\cdots&\mathbf{x}_{n}\end{bmatrix}^{\top}$, this problem is traditionally framed as best subset selection (BSS) [5, 34], which seeks to minimize $\|\mathbf{y}-\mathbf{X}\boldsymbol{\beta}\|^{2}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\mathbf{x}_{i}^{\top}\boldsymbol{\beta})^{2}$ subject to $\|\boldsymbol{\beta}\|_{\ell_{0}}\leq k$, where $\|\boldsymbol{\beta}\|_{\ell_{0}}$ measures the number of nonzero entries in $\boldsymbol{\beta}$. Since exactly solving the BSS problem is computationally infeasible, early approaches to sparse linear regression comprised greedy search (GS) algorithms, such as forward selection (FS) [21], matching pursuit (MP) [61], and orthogonal matching pursuit (OMP) [19]. These algorithms iteratively select features that maximize the immediate reduction in mean squared error or a proxy thereof. In recent years, however, greedy methods have been largely supplanted by the lasso [50], elastic net [62], and other convex relaxation techniques [25, 33]. This shift stems from the fact that, compared to greedy methods, convex relaxation approaches typically offer stronger theoretical guarantees under milder assumptions [51, 14, 54, 12, 33] and often demonstrate superior empirical performance [50, 62, 31, 12].
Breiman hypothesized in 1996 [8] that greedy methods may suffer from instability [58]—small changes in $\mathbf{y}$ or $\mathbf{X}$ can cause drastic differences in selected features and resulting models—leading to high variance in cross-validation performance estimates and making hyperparameter tuning especially challenging. To address this issue, he proposed stabilizing greedy methods through randomized ensembling, which involves averaging the predictions from multiple estimators, each obtained by applying a greedy method to a perturbed version of the dataset. He investigated perturbation techniques such as bootstrap resampling (“bagging”) and adding noise to the response (“smearing”), and showed empirically, as well as to some extent theoretically, that they led to better stability and prediction performance [7, 8, 9]. However, these randomized ensemble methods did not seem to outperform convex relaxation approaches and were further limited by their high computational cost. They have hence not gained popularity in a sparse linear regression context. On the other hand, when adapted to the setting of nonlinear estimators, such as via random forests (RFs), these ensembling techniques have proved to be wildly successful [10, 15, 26, 44]. The exact reasons for their success remains somewhat elusive and an active topic of research [13, 55, 42, 36, 47, 46, 45, 22, 18, 39, 40, 48].
@@
We argue that the reasons for the predictive effectiveness of RGS go beyond Breiman’s earlier hypothesis of stabilization [8] or his later explanation of variance reduction [10]. Instead, RGS, and randomized ensembling more broadly, serves to enlarge the model space being searched over, thereby shifting the entire bias-variance trade-off curve. Indeed, for a fixed number of greedy search iterations $k$, we show empirically that RGS often has a smaller training error compared to its non-ensembled version (GS), while simultaneously having smaller degrees of freedom, which is a measure of the gap between training and test error. We are able to make this argument even more precise through theoretical analysis. Following a longstanding precedent in statistics, we focus on the case of orthogonal features, under which FS, MP, and OMP are all equivalent to BSS. In this setting, we show RGS adaptively reweights ordinary least squares (OLS) coefficients, creating a smooth approximation to the hard thresholding utilized by BSS. These weights are well-approximated by a logistic function in the OLS coefficient rankings, with inflection point roughly equal to $k$ and growth rate depending on the feature subsampling ratio $\gamma\coloneqq m/p$. Varying $\gamma$ for a fixed $k$ thereby searches over a larger space of models compared to BSS. Our first main theoretical result is a lower bound on the resulting decrease in training error in terms of the OLS coefficients. Meanwhile, our second main result identifies a sufficient concavity condition on the OLS coefficients for when RGS is guaranteed to have fewer degrees of freedom compared to BSS for the same value of $k$.
The rest of this paper is organized as follows. We describe RGS and its efficient implementation in Section 2. We present our main theoretical results regarding the reduction in training error and degrees of freedom from feature subsampling in Section 3 and our numerical experiments in Section 5. Further results on the properties of RGS weights are shown in Section 4. In Section 6, we discuss how our theoretical results add clarity to the debate regarding the role of randomization in the success of RFs [55, 42, 18, 39, 40, 48]. Specifically, while previous work argued that feature subsampling works by “regularizing” RFs, our work supports the claim that it can also reduce bias by enlarging the model search space. In Section 7, we conclude with a discussion of the limitations of our study and potential future work.
2 Randomized Greedy Search
In this section, we formally introduce our proposed method, RGS, for solving sparse linear regression. Note that such a method has generality beyond linear modeling, as one can make use of a dictionary of derived features, which could consist of, for example, trigonometric, spline, wavelet, radial, or ridge functions. Such features can either be defined using prior knowledge, or even learned from data. For conceptual simplicity, however, we use the former framing for the rest of this paper. The inner product between two vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ is defined as $\langle\mathbf{a},\mathbf{b}\rangle=(\mathbf{a}^{\top}\mathbf{b})/n$, with norm $\|\mathbf{a}\|^{2}=(\mathbf{a}^{\top}\mathbf{a})/n$. We denote the $j$-th feature vector in bold as $\mathbf{x}_{\cdot j,p}\coloneqq(x_{1j},x_{2j},\dots,x_{nj})^{\top}$ and assume it is normalized so that $\|\mathbf{x}_{\cdot j,p}\|=1$. Functions, such as the $j$-th coordinate function $x_{j,p}$, will be denoted in regular font.
2.1 Forward Selection
We first describe FS, which is the default base learner used by RGS. Starting with an empty active set $\mathcal{M}_{0}=\emptyset$ and iterating over $l=1,2,\ldots,k$ for some prespecified sparsity level $k$, it updates $\mathcal{M}_{l-1}$ with the feature indexed by
$j_{l}$ $\coloneqq\operatorname*{arg\,min}_{j\in\mathcal{V}_{l}}\|\mathbf{y}-\mathbf{P}_{\mathcal{M}_{l-1}\cup\{j\}}\mathbf{y}\|$ (1)
$=\operatorname*{arg\,max}_{j\in\mathcal{V}_{l}}\frac{|\langle\mathbf{r}_{l-1},\mathbf{x}_{\cdot j,p}\rangle|}{\|\mathbf{P}^{\perp}_{\mathcal{M}_{l}}\mathbf{x}_{\cdot j,p}\|},$
and sets the updated model predictions to be $\hat{\mathbf{f}}_{l,p}\coloneqq\mathbf{P}_{\mathcal{M}_{l}}\mathbf{y}$. Here, $\mathbf{P}_{\mathcal{M}_{l}}$ (resp. $\mathbf{P}^{\perp}_{\mathcal{M}_{l}}$) denotes the projection matrix onto the (resp. orthogonal complement of the) column span of the feature vectors indexed in $\mathcal{M}_{l}$, $\mathbf{r}_{l-1}\coloneqq\mathbf{y}-\hat{\mathbf{f}}_{l-1,p}$ denotes the vector of residuals at the $(l-1)$-th step, and $\mathcal{V}_{l}=[p]\backslash\mathcal{M}_{l-1}$. We use $\hat{f}_{k,p}$ to denote the predictive function obtained after $k$ steps, henceforth called the FS estimator.
FS is of course not the only greedy search algorithm. Notable alternatives include OMP, which does not project out the contributions from $\mathbf{x}_{\cdot j,p}$ in (1) and instead just uses the absolute correlation of $\mathbf{r}_{l-1}$ with $\mathbf{x}_{\cdot j,p}$, and MP, which uses the same objective as OMP, but updates the model predictions via $\hat{\mathbf{f}}_{l,p}=\hat{\mathbf{f}}_{l-1,p}+\langle\mathbf{r}_{l-1},\mathbf{x}_{\cdot j,p}\rangle\mathbf{x}_{\cdot j,p}$ and allows the same feature to be selected multiple times. When efficiently implemented, all three algorithms have similar computational complexity. In particular, FS can be implemented by maintaining a QR factorization of the active submatrix $\mathbf{X}_{\mathcal{M}_{l}}$, which can be updated with a rank-one modification upon adding a new selected variable. At step $l$, the computational cost is $O(np)$ for computing (1), $O(nl)$ for updating the QR factorization, and $O(n)$ for updating the residuals. Summing over $k$ steps gives a total cost of $O(npk)$. Since $k\leq\min\{n,p\}$, this cost is no more than that of a single OLS fit.
2.2 Randomized Greedy Search
Our proposed RGS estimator comprises an ensemble of randomized FS estimators. [^2] We call our method RGS because the base learner can be easily replaced with other GS methods, and can be further generalized to GS methods for nonlinear models. Each FS estimator is randomized at each step $l$ by modifying (1) to instead select the maximizer over $j\in\mathcal{V}_{l}$, where $\mathcal{V}_{l}$, called the candidate set, is now a subset of size $m$ drawn uniformly at random from $[p]\backslash\mathcal{M}_{l-1}$. [^3] If $|[p]\backslash\mathcal{M}_{l-1}|<m$, we do not randomize. Stopping at $k$ steps as before, we call the resulting estimator $\tilde{f}_{k,p,b}$. We repeat this model building process across $B$ independent replicates, and aggregate the resulting estimators via simple averaging to get the RGS estimator $\hat{f}_{k,m,p,B}\coloneqq B^{-1}\sum_{b=1}^{B}\tilde{f}_{k,p,b}$. Recall that we call this randomization scheme feature subsampling, and note that it differs from bagging and smearing in that the randomization is interleaved with optimization via the modified (1), whereas bagging and smearing make use of a one-off dataset randomization prior to running FS.
2.3 Efficient RGS via Dynamic Programming
If computed naively, the computational cost of RGS scales linearly in $B$, the number of replicates. Since it is standard to take $B$ to be in the hundreds (or thousands) in order to minimize the contribution of algorithmic variance to prediction error, this could potentially make RGS prohibitively expensive. Fortunately, we are able to design an efficient dynamic programming implementation of RGS that significantly reduces the computational overhead.
Our method leverages the observation that the same active set may appear across multiple replicates, which means that much computation is needlessly repeated. To avoid this waste, we run all replicates simultaneously, maintaining, at the start of each step $l$, an object $\mathcal{S}_{l-1}$, comprising all unique subsets of size $l-1$ selected at the previous step, as well as the number of times each subset occurs. To update $\mathcal{S}_{l-1}$, we process each entry $(\mathcal{M},B_{\mathcal{M}})$ as follows: We generate $B$ candidate sets $\mathcal{V}^{(1)},\mathcal{V}^{(2)},\ldots,\mathcal{V}^{(B)}\subset[p]\backslash\mathcal{M}$ uniformly at random. To solve (1) for all candidate sets, we need only compute the correlations of the residual with features in $\cup_{b=1}^{B}\mathcal{V}^{(b)}$, and likewise for the feature projections, just once in total. Updating $\mathcal{M}$ with these results provides $B$ new, possibly repeated active sets $\mathcal{M}^{(1)},\mathcal{M}^{(2)},\ldots,\mathcal{M}^{(B)}$, which we then return as a list (see also Figure 1). After processing all entries in $\mathcal{S}_{l-1}$ in this manner, we concatenate all the resulting lists and count frequencies to get $\mathcal{S}_{l}$. After $k$ steps, the RGS estimator can then be recovered as $\hat{f}_{k,m,p,B}=B^{-1}\sum_{(\mathcal{M},B_{\mathcal{M}})\in\mathcal{S}_{k}}B_{\mathcal{M}}\tilde{f}_{\mathcal{M}}$, where $\tilde{f}_{\mathcal{M}}$ is the OLS estimator given the data matrix $\mathbf{X}_{\mathcal{M}}$.
For $l=1,2,\ldots,k$, let $\phi(l,m,B)$ denote the expected cardinality of $\mathcal{S}_{l}$. Arguing similarly to FS, the computational cost for handling each active set is $O(np)$. The total cost for our RGS implementation is thus $O\big{(}np\sum_{l=1}^{k}\phi(l,m,B)\big{)}$, which can be much smaller than the naive $O(nmkB)$. We may further reduce the computational cost by repeatedly resampling $\mathcal{S}_{l}$ before processing it. That is, we treat $\mathcal{S}_{l}$ as a sample of size $B$ from a population whose elements are subsets and repeatedly replace $\mathcal{S}_{l}$ with a bootstrap resample of itself. Similar to the regular bootstrap, this progressively reduces the number of distinct elements in $\mathcal{S}_{l}$, while preserving the property that $\mathcal{S}_{l}$ is an unbiased estimate of the original.
Remark.
It is easy to see that this efficient implementation is unique to RGS, and is unavailable to other ensemble methods, which rely on perturbing data, and tend not to produce repeated calculations across replicates. It is also for this reason that we do not combine feature subsampling with bootstrapping, as practiced in RFs. We also note that there are limits to the efficiency gain in edge case hyperparameter choices (i.e., very small $m$ or large $k$).
Figure 1: Diagram showing how the dynamic programming implementation of RGS processes an active set $\mathcal{M}$. It first uses (1) to compute scores for all columns of $\mathbf{X}$ not in $\mathcal{M}$. Simultaneously, it generates $B$ candidate subsets $\mathcal{V}^{(1)},\mathcal{V}^{(2)},\ldots,\mathcal{V}^{(B)}$ of size $m$ from the complement of $\mathcal{M}$. For within each candidate set, it then computes the index with the maximum score. $\mathcal{M}$ is then updated with each of these values, returning a collection $\mathcal{M}^{\prime(1)},\ldots,\mathcal{M}^{\prime(B)}$.
2.4 Prior Research on Feature Subsampling
The idea of combining models that are randomized via feature subsampling was first introduced by Breiman in the context of RFs [10]. It has since been applied in various statistical problems, such as linear regression [36], M-estimation [56], feature importance uncertainty estimation [28], clustering [27], and graph learning [57]. Most closely related to ours are two works that analyze feature subsampling applied to greedy algorithms, which we describe in more detail.
Mentch and Zhou [42] introduce a version of RGS that incorporates bagging. However, their primary aim was to use it as a proxy to study RFs, and did not intend their method to be of practical use. Notably, they did not mention the efficient dynamic programming strategy that we introduced, nor did they provide any theoretical analysis. We will explain their arguments further in Section 6 when we discuss connections of our work to RFs. Separately, Zhan, Zhang, and Xia [59] propose ensemble projection pursuit regression, which applies feature subsampling to projection pursuit, a variant of greedy search that is essentially OMP with ridge-function dictionaries, without using bagging. While they obtain consistency rates for their method, they do not discuss the differences between the ensemble and non-ensemble versions. In particular, they do not show how randomized ensembling can shift the bias-variance trade-off curve.
3 Provable Error Reduction
3.1 Analytical Framework
In this section, we provide a theoretical explanation of the role of randomization in RGS. To do so, we assume a generative model $y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i}$, where $f(\mathbf{x})=\boldsymbol{\beta}^{\top}\mathbf{x}$ is a linear function, $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are IID noise variables with zero mean and variance $\sigma^{2}$ and that are independent of the data matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$. [^4] Strictly speaking, we do not need to assume a well-specified linear model for our main results to hold, but we do so here for simplicity.
Our central claim is that randomization can simultaneously increase model fit and reduce model complexity, in contrast to earlier explanations. In a fixed design setting, the complexity of an estimator $\hat{f}$ is often quantified via its degrees of freedom [31]:
$\operatorname{df}(\hat{f})\coloneqq\sum_{i=1}^{n}\frac{\operatorname{Cov}(y_{i},\hat{f}(\mathbf{x}_{i}))}{\sigma^{2}}.$ (2)
Indeed, letting $\mathbf{f}=\mathbf{X}\boldsymbol{\beta}$ and $\hat{\mathbf{f}}=\mathbf{X}\hat{\boldsymbol{\beta}}$ denote the true and fitted function values respectively, the expected mean squared prediction error of $\hat{f}$ can be decomposed into its expected training error and a constant multiple of its degrees of freedom; that is,
$\mathbb{E}\big{[}\|\mathbf{f}-\hat{\mathbf{f}}\|^{2}\big{]}+\sigma^{2}=\mathbb{E}\big{[}\|\mathbf{y}-\hat{\mathbf{f}}\|^{2}\big{]}+\frac{2\sigma^{2}}{n}\operatorname{df}(\hat{f}).$ (3)
@@ 3.1 Analytical Framework
Therefore, we will analyze how RGS compares to FS on each term on the right hand side of (3).
To aid our analysis, we make some simplifying assumptions. First, we study the performance of the method in the limit as $B\rightarrow\infty$. By the strong law of large numbers, the estimator $\hat{f}_{k,m,p,B}$ converges almost surely to its expected value, with the expectation taken over the randomness of the subset choice. For the rest of our theoretical discussion, we will use and refer to this large-limit model as the RGS estimator, denoting it as $\hat{f}_{k,m,p}$. In addition, we assume that the features are orthogonal:

### v3

@@ Abstract
Feature subsampling is a core component of random forests and other ensemble methods. While recent theory suggests that this randomization acts solely as a variance reduction mechanism analogous to ridge regularization, these results largely rely on base learners optimized via ordinary least squares. We investigate the effects of feature subsampling on greedy forward selection, a model that better captures the adaptive nature of decision trees. Assuming an orthogonal design, we prove that ensembling with feature subsampling can reduce both bias and variance, contrasting with the pure variance reduction of convex base learners. More precisely, we show that both the training error and degrees of freedom can be non-monotonic in the subsampling rate, breaking the analogy with standard shrinkage methods like the lasso or ridge regression. Furthermore, we characterize the exact asymptotic behavior of the estimator, showing that it adaptively reweights OLS coefficients based on their rank, with weights that are well-approximated by a logistic function. These results elucidate the distinct role of algorithmic randomization when interleaved with greedy optimization.
Keywords: Ensemble Learning; Forward Selection; Greedy Model Search; Random Forest; Regression
@@ 1 Introduction
Introduced by breiman2001random, random forests (RFs) are ensembles of decision trees, each grown via greedy, recursive partitioning of the input space (breiman1984classification). Each tree is independently randomized in two ways. First, it is trained on a bootstrap sample of the data. Second, at every internal node it considers only a randomly drawn subset of $m$ (out of $p$ total) features when searching for the next split.
This combination of aggregation and randomized tree construction has proven remarkably effective, making RFs among the most widely used machine learning methods today. Their performance is often close to state-of-the-art on tabular data (caruana2006empirical; fernandez2014we; olson2018data; grinsztajn2022tree), which is prevalent across healthcare, finance, and scientific domains.
There has been significant theoretical work on the consistency of RFs, especially in high dimensions (scornet2015consistency; klusowski2020sparse; syrgkanis2020estimation; chi2022asymptotic; klusowski2024large; mazumder2024convergence), but there is not yet a complete consensus on how randomization improves prediction performance. breiman2001random originally explained the benefits of randomization as increasing diversity among the trees, thereby reducing the variance in their predictions. Interestingly, in his concluding remarks, he speculated that randomization may “act to reduce bias” but that “the mechanism for this is not obvious”. These two explanations have had divergent fates, with hastie2009elements arguing in their influential textbook that “the improvements in prediction obtained by bagging or random forests are solely a result of variance reduction”.
Seminal work by mentch2020randomization elaborated on this “pure variance reduction” hypothesis by offering a quantitative analysis of the role played by the feature subsampling ratio $\gamma=m/p$. In their numerical experiments, they found that decreasing $\gamma$ decreases the degrees of freedom in RFs. To further explore the generality of this phenomenon, they applied ensembling with feature subsampling to the greedy forward selection (FS) algorithm for sparse linear regression (draper1966applied), finding that decreasing $\gamma$ also decreases the degrees of freedom of the resulting family of estimators. Finally, in concurrent work with lejeune2020implicit, they analyzed ensembling with feature subsampling on ordinary least squares (OLS) estimators. Such an ensemble remains a linear estimator, unlike ensembles of FS. Assuming orthonormal features, they found that the resulting regression coefficients were shrunk from the OLS values by a factor of $\gamma$, making the algorithm equivalent to ridge regression. They hence concluded that feature subsampling “serves as a form of implicit regularization” and that $\gamma$ “serves much the same purpose as the shrinkage penalty in explicitly regularized regression procedures like lasso and ridge regression”.
Recent work building on mentch2020randomization has proceeded in two directions. In one direction, patil2024asymptotically relaxed the orthonormal feature assumption to one they called “asymptotic freeness” (relative to the feature subsampling operator). Under this assumption, they showed that ensembling OLS estimators with column-wise sampling (“feature subsampling”) is asymptotically equivalent to ridge regression, albeit with inflated variance. Meanwhile, other works studying row-wise sampling (“data subsampling”) established a broad theoretical equivalence between this form of randomized ensembling and ridge penalization (du2023subsample; patil2023generalized). Together, these works seem to strengthen the analogy between various forms of subsampling randomization and ridge-like shrinkage.
In the other direction, studies by curth2024random, liu2025randomization, mei2024exogenous, and revelas2025random suggest that feature subsampling can also reduce bias in RFs via two mechanisms. The first is enlarging the effective model class (curth2024random). The second is overcoming failures of greedy optimization (liu2025randomization). These results challenge the “pure variance reduction” hypothesis of hastie2009elements and suggest that the shrinkage analogy in mentch2020randomization seems appropriate when describing ensembles of base learners fit with convex optimization, but this analogy becomes strained when subsampling is interleaved with greedy optimization, as is the case for RFs. Existing theory on subsampling-based ensembling primarily concerns non-adaptive or analytically tractable base learners arising from convex optimization, or oracle versions of RFs (mei2024exogenous). For settings in which feature subsampling is interleaved with greedy, adaptive fitting, evidence for bias-reduction mechanisms remains largely empirical.
The goal of our paper is to theoretically investigate the nuances of ensembling with feature subsampling that arise in this more complex scenario. Inspired by mentch2020randomization, we study their ensembling strategy as applied to FS rather than RFs. This model captures the key adaptive, greedy behavior of RFs and allows for rigorous mathematical analysis that would otherwise be intractable. We assume a fixed design setting, under which the expected training error captures the contribution of bias and the degrees of freedom quantify the estimator’s sensitivity to noise and thus its variance component. We further assume orthogonal features, in which case FS is equivalent to best subset selection (BSS, hastie2020extended). Letting FS($k$) and EFS($k,m$) refer to $k$-step forward selection and its ensemble, respectively, we obtain the following theoretical results.
  • For $p$ large enough, the training error of EFS($k,m$) is non-monotonic in $m$, with a minimum achieved at an intermediate value.
  • If the degrees of freedom of FS($k$) forms a concave sequence in $k$, then the degrees of freedom of EFS($k,m$) is increasing in $m$.
  • If the concavity condition does not hold, then the degrees of freedom of EFS($k,m$) may be non-monotonic in $m$, with a minimum achieved at an intermediate value.
These results provide rigorous proof that feature subsampling randomization can simultaneously reduce both bias and variance. On the other hand, variance reduction is not guaranteed but depends on certain structural assumptions. Furthermore, in both ridge and lasso regression, the expected training error (resp. degrees of freedom) is monotonically increasing (resp. monotonically decreasing) in the regularization parameter $\lambda$ (zou2007degrees; hastie2009elements), yielding an exact trade-off between the two quantities. Since this monotonic relationship does not hold for $m$ in EFS($k,m$), we see that shrinkage is not quite an appropriate analogy for the role of randomization in this setting.
In addition, focusing on the asymptotic regime $m,p\to\infty$ with $m/p\to\gamma$, we characterize the exact effect that $\gamma$ has on the regression coefficients of EFS($k,m$). Note that it is standard practice to take $m=\lfloor p/3\rfloor$ in RFs, corresponding to $\gamma=1/3$. Specifically, we show that EFS($k,m$) adaptively reweights OLS coefficients, creating a smooth approximation to the hard thresholding utilized by FS($k$). These weights are well-approximated by a logistic function in the OLS coefficient rankings, with inflection point roughly equal to $k$ and growth rate depending on $\gamma$.
Varying $\gamma$ (equivalently, $m$) for a fixed $k$ can be thought of as reallocating a fixed weight budget among the different features, thereby searching over a larger family of models compared to FS($k$). Since increasing $\gamma$ simultaneously shrinks some coefficients and inflates others, this reveals another difference from ridge and lasso regression, which shrink all coefficients towards zero. Indeed, $\gamma$ has more affinity with the $\alpha$ parameter in elastic net (zou2005regularization), which interpolates between the $\ell_{1}$ and $\ell_{2}$ penalties. However, the family of solutions obtained by EFS($k,m$) is not equivalent to that from elastic net, revealing yet more subtleties in the relationship between explicit regularization and “algorithmic regularization”.
We provide proof sketches for some of our main results. Full details appear in the Supplementary Material (chen2025revisiting-supp).
2 Preliminaries
2.1 Set-up and notation
Assume a generative model $y_{i}=f(\mathbf{x}_{i})+\varepsilon_{i}$, where $f(\mathbf{x})=\boldsymbol{\beta}^{\top}\mathbf{x}$ is a linear function and $\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{n}$ are IID noise variables with zero mean and variance $\sigma^{2}$. Strictly speaking, we do not need to assume a well-specified linear model for our main results to hold, but we do so here for conceptual simplicity. We assume $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{n}]^{\top}\in\mathbb{R}^{n\times p}$ is a fixed design matrix. For vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$, we define the inner product as $\langle\mathbf{a},\mathbf{b}\rangle=\mathbf{a}^{\top}\mathbf{b}/n$, with associated norm $\|\mathbf{a}\|^{2}=\langle\mathbf{a},\mathbf{a}\rangle$. For each feature index $j\in[p]$, let $\mathbf{x}_{\cdot j,p}=(x_{1j},x_{2j},\dots,x_{nj})^{\top}$ denote the $j$-th feature vector, which we assume to be normalized so that $\|\mathbf{x}_{\cdot j,p}\|=1$. We use regular font to denote functions such as the coordinate functions $x_{j,p}$.
2.2 Algorithms
Forward selection.
Starting with an empty active set $\mathcal{M}_{0}=\emptyset$, forward selection (FS) proceeds for $l=1,2,\dots,k$ by updating $\mathcal{M}_{l-1}$ with the index
$j_{l}=\operatorname*{arg\,min}_{j\in\mathcal{V}_{l}}\|\mathbf{y}-\mathbf{P}_{\mathcal{M}_{l-1}\cup\{j\}}\mathbf{y}\|=\operatorname*{arg\,max}_{j\in\mathcal{V}_{l}}\frac{|\langle\mathbf{r}_{l-1},\mathbf{x}_{\cdot j,p}\rangle|}{\|\mathbf{P}_{\mathcal{M}_{l-1}}^{\perp}\mathbf{x}_{\cdot j,p}\|},$ (1)
where $\mathcal{V}_{l}=[p]\setminus\mathcal{M}_{l-1}$ is the set of eligible indices. Here, $\mathbf{P}_{\mathcal{M}_{l}}$ is the projection matrix onto the column span of the selected features, $\mathbf{P}_{\mathcal{M}_{l}}^{\perp}$ projects onto its orthogonal complement, and $\mathbf{r}_{l-1}=\mathbf{y}-\hat{\mathbf{f}}_{l-1,p}$ denotes the residual vector at step l1l-1. We set
$\hat{\mathbf{f}}_{l,p}=\mathbf{P}_{\mathcal{M}_{l}}\mathbf{y}.$
After $k$ steps, the resulting predictive function $\hat{f}_{k,p}$ is referred to as the FS($k$) estimator.
Ensemble forward selection.
Following mentch2020randomization, we introduce randomization into each FS($k$) step by modifying (1) so that the maximization is taken over a randomly drawn subset. Specifically, at step $l$, we sample a candidate set $\mathcal{V}_{l}\subseteq[p]\setminus\mathcal{M}_{l-1}$ of size $m$ uniformly without replacement. If $|[p]\setminus\mathcal{M}_{l-1}|<m$, we use all remaining features without randomization. The update index $j_{l}$ is then chosen as in (1), but restricted to $j\in\mathcal{V}_{l}$. Stopping after $k$ steps yields a randomized FS($k$) estimator $\tilde{f}_{k,p,b}$ for replicate $b$. Repeating this process independently for $b=1,\dots,B$ and aggregating via simple averaging yields the ensemble estimator
$\hat{f}_{k,m,p,B}=\frac{1}{B}\sum_{b=1}^{B}\tilde{f}_{k,p,b}.$
By the strong law of large numbers, the estimator $\hat{f}_{k,m,p,B}$ converges almost surely to its expected value as $B\to\infty$, with the expectation taken over the randomness of the subset choices. For the rest of this paper, we use and refer to this limit model as the EFS($k,m$) estimator, denoting it as $\hat{f}_{k,m,p}$.
2.3 Prediction error decomposition
In a fixed design setting, the complexity of an estimator $\hat{f}$ is often quantified via its degrees of freedom (hastie2009elements),
$\operatorname{df}(\hat{f})=\sum_{i=1}^{n}\frac{\operatorname{Cov}(y_{i},\hat{f}(\mathbf{x}_{i}))}{\sigma^{2}}.$ (2)
Letting $\mathbf{f}=\mathbf{X}\boldsymbol{\beta}$ and $\hat{\mathbf{f}}=\mathbf{X}\hat{\boldsymbol{\beta}}$ denote the true and fitted function values respectively, the expected mean squared prediction error of $\hat{f}$ can be decomposed into its expected training error and a constant multiple of its degrees of freedom,
@@ 2.3 Prediction error decomposition
We will therefore analyze how EFS($k,m$) compares to FS($k$) on each term on the right-hand side of (3).
3 Properties of EFS($k,m$) weights
3.1 EFS($k,m$) as reweighted OLS
For the rest of this paper, we assume that the features are orthogonal.